what is the probability that a market inspector will discover greater three violations of the public health code



1
Expert's answer
2022-02-21T18:56:21-0500

The probability that a market inspector will discover more than three violations of the public health code is"P(X \gt 3) = (1-p(X\le3))=1-\{P(X=0)+P(X=1)+P(X=2)+P(X=3)\}"